OpenStudy (anonymous):

(4x + 3y) - (2x -5y) (Hint: Remember to distribute the negative to the second set of parentheses.)

OpenStudy (anonymous):

4-2 = ? 3+5 = ?

OpenStudy (anonymous):

2 8

OpenStudy (turingtest):

yes, do you know the answer now?

OpenStudy (anonymous):

i think so

OpenStudy (turingtest):

what do you have?

OpenStudy (anonymous):

what you mean

OpenStudy (turingtest):

what did you get for your answer?

OpenStudy (anonymous):

2 and 8

OpenStudy (turingtest):

No, that is not even the right form of the answer. What happened to the x and y? I don't think you understand the question as well as henkjan thought you did. first of all, do you know how to simplify -(2x-5y) ??? as it says above, you need to "distribute the negative sign" do you know what that means? what do you get by doing that?

OpenStudy (anonymous):

um does it stay like that

OpenStudy (turingtest):

no, the rule of distribution works like this in general:\[a(x+y)=ax+ay\]for your problem we treat "a" as -1, so we have\[-1*(2x-5y)=-2x+5y\](remember that minus times minus is positive)

OpenStudy (anonymous):

ok thank you

OpenStudy (turingtest):

I hope you do not think that is the answer above...

OpenStudy (anonymous):

no lol

OpenStudy (turingtest):

Do you just not care to continue, or do you have the answer? If you do, what is it?

OpenStudy (anonymous):

3xy